OpenStudy (anonymous):

Solve 6^x = 24 a. x ≈ 30.84 b. x ≈ 1.77 c. x ≈ 0.56 d. x ≈ 0.23

OpenStudy (jackellyn):

Do the natural log (or log) of both sides ln(6^x)=ln 24 x(ln 6) = ln 24 Divide both sides by (ln 6) x= ln 24/ ln 6

OpenStudy (anonymous):

I see what you guys are doing but these are the answers I was given...

OpenStudy (jackellyn):

the answer I gave you, if you put in your calculator x=ln 24/ln 6 it will give you one of the choices. can you do the calculation and say which one it is?

OpenStudy (anonymous):

yes! i got 1.79, but the little almost equal sign means it's b. right?

OpenStudy (jackellyn):

that is right @J_Kerry2! That sign means "about" so it is about 1.77

OpenStudy (jackellyn):

@sabimaru I am not sure there is a rule for what you did. ln 18 = 2.89037176 while the answer was ~1.77.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

@jackellyn - typed this quickly and miffed...was thinking of ln(a/b) = ln(a)-ln(b). Sorry for the confusion!

OpenStudy (jackellyn):

@sabimaru no problem! i was just making sure you knew :)